# Assessing the Reasoning Capabilities of LLMs in the context of Evidence-based Claim Verification

### v3

@@ Abstract
John Dougrez-Lewis1,∗  Mahmud Elahi Akhter2,∗  Federico Ruggeri3 Sebastian Löbbers2Yulan He4,5Maria Liakata2,5 University of Warwick, UK1 Queen Mary University of London, UK2 University of Bologna, Italy3 King’s College London, UK4 The Alan Turing Institute, UK5 j.dougrez-lewis@warwick.ac.uk, yulan.he@kcl.ac.uk, federico.ruggeri6@unibo.it {m.akhter, s.lobbers, m.liakata}@qmul.ac.uk These authors contributed equally to this work.
@@ 1 Introduction
These findings motivate the need for further research on the reasoning capabilities of LLMs, especially in high-stake real-world applications, where research on this topic is in its infancy. A notable example is fact-checking, where LLMs are considered to hold great potential for increased productivity overshadowed by the ease with which bad actors can proliferate misinformation Guo et al. (2023). Verifying information is challenging since models require both accurate classification and strong rationale generation to be effective Schlichtkrull et al. (2023). It is thus essential to understand the reasoning capabilities and limitations of LLMs in the context of fact-checking. In particular, we extend the current discussion around the reasoning abilities of LLMs, focusing on their ability to verify real-world claims.
@@
In this work, we first propose a framework for breaking down complex claims into atomic reasoning steps. The motivation behind this is the lack of uniform terminology around reasoning evaluation. Most prominent evaluation datasets for reasoning are based on mathematics and coding, which involves deductive reasoning but is not treated as such  Sprague et al. (2024). Our framework is rooted in existing philosophy literature concerning logical reasoning that aligns well with NLP Wason and Johnson-Laird (1972); Galotti (1989).
@@
We use our framework to create Reasoning in Evidence-based Claim Verification (RECV), the first reasoning benchmark for claim-verification. The benchmark comprises three datasets, curated from existing resources targeting different domains: VitaminC Schuster et al. (2021) from Wikipedia, CLIMATE-FEVER Diggelmann et al. (2020) from online claims and Wikipedia, and PHEMEPlus Dougrez-Lewis et al. (2022) from rumours circulating on social media and associated evidence from news articles. The claims involve increasing levels of complexity as we move from VitaminC to PHEMEPlus, often requiring deductive and/or abductive reasoning.
@@
We use RECV to evaluate three state-of-the-art proprietary LLMs that have shown impressive performance on various reasoning and language benchmarks Huang and Chang (2023); DeepSeek-AI et al. (2025). These models are: Claude V3 Sonnet Anthropic (2023), GPT-4 OpenAI (2023), and GPT-4o OpenAI et al. (2024). In particular, we prompt models with and without Chain-of-Thought (CoT) Wei et al. (2023) rationale generation to assess if and how the latter influences reasoning. In alignment with previous work Saparov et al. (2023); Akyürek et al. (2024); Li et al. (2024) we find that LLMs are capable of deductive reasoning. However, they consistently fail at claim verification when presented with evidence that requires abductive reasoning. Furthermore, we observe conflicting results when prompting LLMs with CoT strategies. In particular, CoT leads to performance improvements for simple claim verification as in VitaminC, but detrimental in the case of complex claims such as those found in CLIMATE-FEVER and PHEMEPlus. Lastly, we carry out a qualitative analysis of generated rationales and observe high semantic similarity with human explanations, especially in deductive reasoning cases.
In summary, we make the following contributions:
@@
    We show that generated rationales are consistent with human reasoning for correct predictions, but the model is often unable to leverage such rationales for claim verification (§6).
@@ 2 RECV Logical Framework
In particular we consider that reasoning consists of the interplay of three interrelated components: types, processes, and tasks. This is the basis of our RECV framework. Reasoning types are different forms of logical inference that we can use to reach a conclusion from a set of observations or premises. We distinguish between atomic and compound reasoning types. Atomic types denote basic forms of logical inference and include deductive, abductive, inductive, and analogical reasoning.
A reasoning task is any task that requires multiple reasoning types, often in complex interaction with each other. For example claim verification is a composite reasoning task. A reasoning process is the method of interaction between reasoning types or even tasks within complex reasoning tasks. Notable examples of reasoning processes are multi-hop or multi-step inference, where individual steps or hops can be of different reasoning types. In this paper we focus particularly on atomic reasoning types.
@@ 3 Methodology
The application of the RECV framework can be seen in Figure 1. The reasoning task here is claim verification and the reasoning type is composite. The claim is resolved using a single-step process, that consists of abductive type atomic reasoning. Here we only highlight the most plausible hypothesis that resolves the claim as true. However, in practice, we would generate multiple hypotheses before coming to the most plausible one.
@@
Our objective here was to determine the atomic reasoning types necessary for accomplishing claim verification from evidence. We first collected a small dataset by manually selecting 90 claims and associated evidence from VitaminC Schuster et al. (2021), CLIMATE-FEVER Diggelmann et al. (2020), and PHEMEPlus Dougrez-Lewis et al. (2022). We focus on these resources as they are widely used in claim and rumour verification and differ in complexity.
Two annotators with expertise in Computer Science and native English proficiency assigned reasoning type labels to claim-evidence pairs following (§2). Disagreements encountered were resolved via a discussion stage with an independent expert. The Inter-Annotator Agreement (IAA) measured as Bennett’s S score Bennet et al. (1954) to account for label imbalance of reasoning types is 0.90, denoting almost perfect agreement. We observe that all examples either require deductive or abductive reasoning types.
@@
Our preliminary investigation suggests that inductive and analogical reasoning are rarely employed in claim verification. This is presumably because inductive reasoning relies on complete evidence, which is rarely available in real-world domain-specific settings. Generalisations from one domain to another, relevant to inductive reasoning, may only occur in scenarios that share common background knowledge, as in the medical domain.
Similarly, analogical reasoning may be more suitable for other fact-checking related tasks like profiling and motive analysis where comparing information bits frequently occurs to reach a conclusion. By contrast, deductive and abductive reasoning types are often required in fact-checking Pan et al. (2023); Tan et al. (2024). For these reasons, here we focus on deduction and abduction. We show that they represent a challenging setting for claim verification (§4), and model evaluation with LLMs (§5).
@@ 4 RECV Benchmark
We build a heuristic-based sampling strategy to mitigate the anticipated data imbalance between deductive and abductive samples, as it was important to ensure both are represented in the annotated data.
We used a combination of three embedding-based text similarity metrics to compute the average claim similarity between deductive samples. Likewise for abductive samples. The metrics are: cosine similarity, BERTScore Zhang et al. (2020) and BLEURT score Pu et al. (2021). We used the data collected during our preliminary study (§3) to set a similarity threshold for each reasoning type. We used each threshold to sample claims likely to be resolved via deductive and abductive reasoning, respectively. In particular, we exclude instances labeled as ‘unverified’ since such claims are always associated with deductive reasoning, either due to lack of proper evidence or contradictory evidence. In total, we sampled 500 claim-evidence pairs from each dataset. See Appendix C for more details on data sampling.
@@
We recruited 9 PhD students in Computer Science, fluent in English and grouped them in triples, one for each dataset. We evenly distributed dataset samples to annotators in a triple, so that 100 samples were annotated by all. In total, each annotator in a triple labeled 233 samples. Annotation guidelines per dataset are in Appendix D. We computed IAA as Bennett’s S score to account for label imbalance (see Appendix E for pairwise agreement scores). The IAA is 0.75 for VitaminC, 0.56 for CLIMATE-FEVER, and 0.67 for PHEMEPlus. Table 1 reports our RECV statistics.
In particular, we observe that the rate of abductive reasoning samples is relatively low compared to deductive ones: 5.8% in VitaminC, 20.4% in CLIMATE-FEVER, and 7.2% in PHEMEPlus. This imbalance is expected given the nature of collected evidence; most evidence provided, either in the form of Wikipedia articles as in VitaminC and CLIMATE-FEVER or news articles as in PHEMEPlus, contains detailed information to deductively verify the claim. In total, RECV consists of 1500 claim-evidence pairs with associated veracity and reasoning labels. The average sentence length for evidence in VitaminC is 1.084, 7.562 for PHEMEPlus and 7.828 for CLIMATE-FEVER. This highlights the varying complexity of the datasets and RECV.
@@ 6 Explanation evaluation
Providing reasonable explanations to support predicted veracity labels is a crucial aspect of claim verification systems. In particular, an automated system needs to be both convincing and trustworthy to convince users in practice Schlichtkrull et al. (2023). Therefore, we evaluate the LLMs generated rationales in the Exp setting. This is paramount considering that LLMs tend to hallucinate Bouyamourn (2023); Rawte et al. (2023) and be self-contradictory at times Mündler et al. (2023). We randomly selected 100 samples from each dataset in RECV and compared generated rationales against those provided by human annotators. We restricted sample selection to those where at least three models predicted wrong veracity labels. We follow Song et al. (2024) and compute Factual Consistency (FC), Evidence Appropriateness (EA), BARTScore Yuan et al. (2021), and Perplexity (PPL) to assess the quality of the generated explanations. We provide additional details about metrics in Appendix G.
@@ 6.1 Results
Table 3 reports the results concerning explanation evaluation. We observe that all models achieve comparable results on appropriateness (EA), consistency (FC), and coherence measured via BARTScore (BART), while showing notable discrepancies regarding perplexity (PPL). In particular, GPT-4o ZS CoT has the most faithful rationales across all datasets. Moreover, prompting strategies like ZS CoT and M-CoT do not lead to consistent improvements over ZS, suggesting that their effectiveness may be problem- and model-dependent.
Additionally, we assess generated rationales regarding correct and wrong model predictions in Appendix G. Our results show that rationales from correct predictions better align with ground-truth explanations, suggesting that wrong predictions are usually the by-product of incorrect reasoning (the model is unable to leverage the explanation).
Lastly, we analyze how similar the generated rationales were between the models. To do so, we perform a permutation test using sentence-level contradiction scores from Fact_Score (see Appendix G). We find that Claude ZS has the most unique rationales on all datasets. We discuss properties of rationales generated in the case of abductive and deductive errors per dataset.
@@ 7 Findings
Claim: Peking University is in a unitary sovereign state that’s located in East Asia.
@@
Our experiments show that prompting strategies like ZS CoT and M-CoT do not lead to systematic performance improvements, but are rather specific to datasets (e.g., VitaminC) and models. These results align with recent findings about CoT being beneficial mainly for math- and code-related tasks (Sprague et al., 2024). This is likely derived by divergent reasoning paths within the models during inference that lead to reduction in performance (Chollet, 2023; Todd et al., 2023; Dutta et al., 2024).
Furthermore, we also observe that internal alignment can hinder reasoning capabilities when it comes to abductive reasoning. Models are averse to provide predictions when evidence is incomplete. Yet abductive reasoning is often required for more complex tasks such as legal reasoning, just in time fact-checking, and other diverse forms of composite reasoning tasks. Hence, in order to achieve good results on these type of reasoning tasks, LLMs need to improve in the direction of abductive reasoning.
@@ 8 Related work
Several contributions have evaluated different reasoning capabilities in LLMs, including atomic and compounds types.For instance, LLMs can perform abductive reasoning for event prediction Shi et al. (2023), but struggle with common sense reasoning Zhao et al. (2023). Similarly, deductive reasoning in LLMs is beneficial to theorem-proving Saparov et al. (2023), factual content generation Akyürek et al. (2024), and question-answering Li et al. (2024). Nonetheless, the observed improvements are often attributable to how prompts are designed rather than an emergent deductive capability Chen et al. (2024). Moreover, LLMs perform out-of-context inductive reasoning Treutlein et al. (2024), but fail in lexical tasks Ye et al. (2023). Regarding analogical reasoning, LLMs address a wide variety of tasks, including nonverbal tests Webb et al. (2023); Hu et al. (2023), question-answering Yu et al. (2023), mathematical problem solving Yasunaga et al. (2024), and planning Yu et al. (2024), but present shortcomings in as many others Ye et al. (2024); Sourati et al. (2024); Stevenson et al. (2024); Ahrabian et al. (2024); Lewis and Mitchell (2024).
Likewise, despite promising results in compound reasoning tasks, such as counterfactual Wu et al. (2023), and compositional reasoning Lu et al. (2023a), LLMs are notably unreliable Gao et al. (2023); Zhang et al. (2024), sensitive to context Hosseini et al. (2024); Chang and Bergen (2024), and rely on shortcuts Yang et al. (2024).
LLMs for Claim Verification
Early work with LLMs focused on verifying simple facts Lee et al. (2020). More recently, LLMs for claim verification have been augmented with external knowledge Li et al. (2023a); Cheung and Lam (2023), prompt-based reasoning Cao (2023); Li et al. (2023b); Lin et al. (2023), claim decomposition for fine-grained search into text chunks Li et al. (2023a) or first-order logic terms Wang and Shu (2023), and data-augmentation Alhindi et al. (2023). While LLMs have been extensively applied in fact-checking, the question of which reasoning capabilities are needed to verify claims remains unaddressed. Thus, we are the first to propose a reasoning benchmark for claim verification.

### v4

@@ Abstract
[…]
@@ 1 Introduction
These findings motivate the need for further research on the reasoning capabilities of LLMs, especially in high-stake real-world applications, where research on this topic is in its infancy. A notable example is fact-checking, where LLMs are considered to hold great potential for increased productivity even if at the same time they also facilitate bad actors in the proliferation of misinformation Guo et al. (2023) Verifying information is challenging since models require both accurate veracity classification and strong rationale generation to be effective Schlichtkrull et al. (2023). It is thus essential to understand the reasoning capabilities and limitations of LLMs in the context of fact-checking. In particular, we extend the current discussion around the reasoning abilities of LLMs, focusing on their ability to verify real-world claims.
@@
In this work, we first propose a framework for breaking down complex claims into atomic reasoning steps. The motivation behind this is the lack of uniform terminology around reasoning evaluation. Most prominent evaluation datasets for reasoning are based on mathematics and coding, which involve deductive reasoning, even though the use of deduction is not made explicit Sprague et al. (2024).
@@
Our framework is rooted in existing philosophy literature concerning logical reasoning that aligns well with NLP Wason and Johnson-Laird (1972); Galotti (1989). We use our framework to create Reasoning in Evidence-based Claim Verification (RECV), the first reasoning benchmark for claim-verification. The benchmark comprises three datasets, curated from existing resources targeting different domains: VitaminC Schuster et al. (2021) from Wikipedia, CLIMATE-FEVER Diggelmann et al. (2020) from online claims and Wikipedia, and PHEMEPlus Dougrez-Lewis et al. (2022) from rumours circulating on social media and associated evidence from news articles. The claims involve increasing levels of complexity as we move from VitaminC to PHEMEPlus, often requiring deductive and/or abductive reasoning.
@@
We use RECV to evaluate three state-of-the-art proprietary LLMs that have shown impressive performance on various reasoning and language benchmarks Huang and Chang (2023); DeepSeek-AI et al. (2025). These models are: Claude V3 Sonnet Anthropic (2023), GPT-4 OpenAI (2023), and GPT-4o OpenAI et al. (2024). In particular, we prompt models with and without Chain-of-Thought (CoT) Wei et al. (2023) rationale generation to assess if and how the latter influences reasoning. In alignment with previous work Saparov et al. (2023); Akyürek et al. (2024); Li et al. (2024) we find that LLMs are capable of deductive reasoning. However, they consistently fail at claim verification when presented with evidence that requires abductive reasoning. Furthermore, we observe conflicting results when prompting LLMs with CoT strategies. In particular, CoT leads to performance improvements for simple claim verification as in VitaminC, but is detrimental in the case of complex claims such as those found in CLIMATE-FEVER and PHEMEPlus. Lastly, we carry out a qualitative analysis of generated rationales and observe high semantic similarity with human explanations, especially in deductive reasoning cases. In summary, we make the following contributions:
@@
    We show that generated rationales are consistent with human reasoning for correct predictions, but model are often unable to leverage such rationales for claim verification (§7).
@@
2 Terminology
To avoid ambiguity, we provide a brief overview of the main terminology used in the paper. We first introduce fact-checking and rumour verification specific terminology. We define a claim as a check-worthy statement, i.e., a piece of information that has to be verified. Similarly, a rumour is a widely circulating claim of unknown veracity Zubiaga et al. (2018). An evidence is a piece of information, here textual, retrieved from a document, that can be used to verify the veracity of a claim.
We also provide an overview of reasoning-specific terminology. We denote conclusion to be a statement that derives from the direct elaboration of some observations. An observation is a piece of factual information. In the case of a reasoning task for claim verification, observations coincide with evidence, as reasoning is performed on factual information with the intent of verifying a given claim. Lastly, we denote explanations as natural language descriptions of the reasoning process that leads to a conclusion from a given set of observations.
@@ 3 RECV Logical Framework
In particular we consider that reasoning consists of the interplay of three interrelated components: types, processes, and tasks. This is the basis of our RECV framework. Reasoning types are different forms of logical inference that we can use to reach a conclusion from a set of observations or premises. We distinguish between atomic and compound reasoning types. Atomic types denote basic forms of logical inference and include deductive, abductive, inductive, and analogical reasoning. A reasoning task is any task that requires multiple reasoning types, often in complex interaction with each other. For example claim verification is a composite reasoning task. A reasoning process is the method of interaction between reasoning types or even tasks within complex reasoning tasks. Notable examples of reasoning processes are multi-hop or multi-step inference, where individual steps or hops can be of different reasoning types. In this paper we focus particularly on atomic reasoning types.
@@ 4 Methodology
The application of the RECV framework can be seen in Figure 1. The reasoning task here is claim verification and the reasoning type is abductive. The claim is resolved using a single-step process, that consists of abductive type atomic reasoning. Here we only highlight the most plausible hypothesis that resolves the claim as true. However, in practice, we would generate multiple hypotheses before coming to the most plausible one.
@@
Our objective here was to determine the atomic reasoning types necessary for accomplishing claim verification from evidence. We first collected a small dataset by manually selecting 90 claims and associated evidence from VitaminC Schuster et al. (2021), CLIMATE-FEVER Diggelmann et al. (2020), and PHEMEPlus Dougrez-Lewis et al. (2022). We focus on these resources as they are widely used in claim and rumour verification and differ in complexity. Two annotators with expertise in Computer Science and native English proficiency assigned reasoning type labels to claim-evidence pairs following (§3). Disagreements encountered were resolved via a discussion stage with an independent expert. The Inter-Annotator Agreement (IAA) measured as Bennett’s S score Bennet et al. (1954) to account for label imbalance of reasoning types is 0.90, denoting almost perfect agreement. We observe that all examples either require deductive or abductive reasoning types.
@@
Our preliminary investigation suggests that inductive and analogical reasoning are rarely employed in claim verification. This is presumably because inductive reasoning relies on complete evidence, which is rarely available in real-world domain-specific settings. Generalisations from one domain to another, relevant to inductive reasoning, may only occur in scenarios that share common background knowledge, as in the medical domain. Similarly, analogical reasoning may be more suitable for other fact-checking related tasks like profiling and motive analysis where frequent and repeated comparisons may occur to reach a conclusion. By contrast, deductive and abductive reasoning types are often required in fact-checking Pan et al. (2023); Tan et al. (2024). For these reasons, here we focus on deduction and abduction. We show that they represent a challenging setting for claim verification (§5), and model evaluation with LLMs (§6).
@@ 5 RECV Benchmark
We build a heuristic-based sampling strategy to mitigate the anticipated data imbalance between deductive and abductive samples, as it was important to ensure both are represented in the annotated data. We used a combination of three embedding-based text similarity metrics to compute the average claim similarity between deductive samples. Likewise for abductive samples. The metrics are: cosine similarity (SimScore), BERTScore Zhang et al. (2020) and BLEURT score Pu et al. (2021). We used the data collected during our preliminary study (§4) to set a similarity threshold for each reasoning type. In particular, we computed the distribution of similarity metrics in the preliminary study data. We computed a separate distribution for deductive and abductive samples, respectively. See Figure A1 in Appendix C for a summary. We found that abductive samples had lower BERTScore and higher lexical similarity (SimScore), likely due to indirect evidence, with low surcace overlap.Deductive samples, by contrast, had higher BLEURT and SimScore. From these observations, we derived two filtering thresholds. For abductive samples, the threshold is: $\text{BERTScore}\leq 0.25\,\wedge\,\text{SimScore}\geq 0.35$. For deductive samples, the threshold is: $\text{SimScore}\geq 0.36\,\wedge\,\text{BLEURT}>0.15$.
We used each threshold to sample claims likely to be resolved via deductive and abductive reasoning, respectively. In particular, we exclude instances labeled as ‘unverified’ since such claims are always associated with deductive reasoning, either due to lack of appropriate evidence or contradictory evidence. In total, we sampled 500 claim-evidence pairs from each dataset. This strategy allowed us to avoid an over-representation of deductive cases,ensuring diversity in the annotated data. However, we remark that this strategy is noisy at best and only provides a weak guarantee. See Appendix C for more details on data sampling.
@@
We recruited 9 PhD students in Computer Science, fluent in English and grouped them in triples, one for each dataset. The annotation task involved labeling claim-evidence pairs as requiring either deductive or abductive reasoning in order to be resolved. More precisely, annotators labeled claim-evidence pairs with the reasoning types deductive or abductive. Table A2 (top) in Appendix D summarizes the annotation process. We evenly distributed dataset samples to annotators in a triple, so that 100 samples were annotated by all. In total, each annotator in a triple labeled 233 samples. Annotation guidelines per dataset are in Appendix D. We computed IAA as Bennett’s S score to account for label imbalance (see Appendix E for pairwise agreement scores). The IAA is 0.75 for VitaminC, 0.56 for CLIMATE-FEVER, and 0.67 for PHEMEPlus. Table 1 reports our RECV statistics. In particular, we observe that the rate of abductive reasoning samples is relatively low compared to deductive ones: 5.8% in VitaminC, 20.4% in CLIMATE-FEVER, and 7.2% in PHEMEPlus. This imbalance is expected given the nature of collected evidence; most evidence provided, either in the form of Wikipedia articles as in VitaminC and CLIMATE-FEVER or news articles as in PHEMEPlus, contains detailed information to deductively verify the claim. In total, RECV consists of 1500 claim-evidence pairs with associated veracity and reasoning labels. The average sentence length for evidence in VitaminC is 1.084, 7.562 for PHEMEPlus and 7.828 for CLIMATE-FEVER. This highlights the varying complexity of the datasets and RECV.
@@ 7 Explanation evaluation
Providing reasonable explanations to support predicted veracity labels is a crucial aspect of claim verification systems. In particular, an automated system needs to be both convincing and trustworthy to convince users in practice Schlichtkrull et al. (2023). Therefore, we evaluate the LLMs generated rationales in the Exp setting. This is crucial considering that LLMs tend to hallucinate Bouyamourn (2023); Rawte et al. (2023) and be self-contradictory at times Mündler et al. (2023). We randomly selected 100 samples from each dataset in RECV and compared generated rationales against those provided by human annotators. We note that the provided human rationales explain the reasons for the chosen veracity and do not explicitly have any mention of why the given rationale corresponds to a reasoning type. In particular, the main goal of this study is to evaluate the models’ reasoning capabilities through implicit measurement and without biasing the models with the mention of reasoning modes. This is also equivalent to how models are prompted, where we do mention reasoning types. We instruct a third annotator to evaluate the quality of provided human rationales. See Appendix G for more details. We restricted sample selection to those where at least three models predicted wrong veracity labels. We follow Song et al. (2024) and compute Factual Consistency (FC), Evidence Appropriateness (EA), BARTScore Yuan et al. (2021), and Perplexity (PPL) to assess the quality of the generated explanations. We provide additional details about metrics in Appendix H.
@@ 7.1 Results
Table 3 reports the results concerning explanation evaluation. Appendix I reports statistical significance results to support our findings. We observe that all models achieve comparable results on appropriateness (EA), consistency (FC), and coherence measured via BARTScore (BART), while showing notable discrepancies regarding perplexity (PPL). In particular, GPT-4o ZS CoT has the most faithful rationales across all datasets. Moreover, prompting strategies like ZS CoT and M-CoT do not lead to consistent improvements over ZS, suggesting that their effectiveness may be problem- and model-dependent.
Additionally, we assess generated rationales regarding correct and wrong model predictions in Appendix H. Our results show that rationales from correct predictions better align with ground-truth explanations, suggesting that wrong predictions are usually the by-product of incorrect reasoning (the model is unable to leverage the explanation).
Lastly, we analyze how similar the generated rationales were between the models. To do so, we perform a permutation test using sentence-level contradiction scores from Fact_Score (see Appendix H). We find that Claude ZS has the most unique rationales on all datasets.
We discuss properties of rationales generated in the case of abductive and deductive errors per dataset.
@@ 8 Findings
Claim: Peking University is a unitary sovereign state that’s located in East Asia.
@@
Our experiments show that prompting strategies like ZS CoT and M-CoT do not lead to systematic performance improvements, but are rather specific to datasets (e.g., VitaminC) and models. These results align with recent findings about CoT being beneficial mainly for math- and code-related tasks (Sprague et al., 2024). This is likely derived by divergent reasoning paths within the models during inference that lead to reduction in performance (Chollet, 2023; Todd et al., 2023; Dutta et al., 2024). Furthermore, we also observe that internal alignment can hinder reasoning capabilities when it comes to abductive reasoning. Models are averse to provide predictions when evidence is incomplete. Yet abductive reasoning is often required for more complex tasks such as legal reasoning, just in time fact-checking, and other diverse forms of composite reasoning tasks. Hence, in order to achieve good results on these type of reasoning tasks, LLMs need to improve in the direction of abductive reasoning.
@@ 9 Related work
Several contributions have evaluated different reasoning capabilities in LLMs, including atomic and compounds types. For instance, LLMs can perform abductive reasoning for event prediction Shi et al. (2023), but struggle with common sense reasoning Zhao et al. (2023). Similarly, deductive reasoning in LLMs is beneficial to theorem-proving Saparov et al. (2023), factual content generation Akyürek et al. (2024), and question-answering Li et al. (2024). He et al. (2025) looked into analogical reasoning along with composite reasoning tasks and came to the conclusion that LLMs were able to handle structured analogical reasoning, but failed at more abstract composite reasoning tasks such as legal reasoning. Nonetheless, the observed improvements are often attributable to how prompts are designed rather than an emergent deductive capability Chen et al. (2024). Moreover, LLMs perform out-of-context inductive reasoning Treutlein et al. (2024), but fail in lexical tasks Ye et al. (2023). Regarding analogical reasoning, LLMs address a wide variety of tasks, including nonverbal tests Webb et al. (2023); Hu et al. (2023a), question-answering Yu et al. (2023), mathematical problem solving Yasunaga et al. (2024), and planning Yu et al. (2024), but present shortcomings in as many others Ye et al. (2024); Sourati et al. (2024); Stevenson et al. (2024); Ahrabian et al. (2024); Lewis and Mitchell (2024). Likewise, despite promising results in compound reasoning tasks, such as counterfactual Wu et al. (2023), and compositional reasoning Lu et al. (2023a), LLMs are notably unreliable Gao et al. (2023); Zhang et al. (2024), sensitive to context Hosseini et al. (2024); Chang and Bergen (2024), and rely on shortcuts Yang et al. (2024).
LLMs for Claim Verification.
Early work with LLMs focused on verifying simple facts Lee et al. (2020). More recently, LLMs for claim verification have been augmented with external knowledge Li et al. (2023a); Cheung and Lam (2023), prompt-based reasoning Cao (2023); Li et al. (2023b); Lin et al. (2023), claim decomposition for fine-grained search into text chunks Li et al. (2023a) or first-order logic terms Wang and Shu (2023), and data-augmentation Alhindi et al. (2023). A close work to ours is Xu et al. (2025). Contrary to Xu et al. (2025), we evaluate models based on real-world claims and evidence. Additionally, in Xu et al. (2025), models are prompted to generate outputs using specific reasoning modes. This likely implicitly biases the models and increases performance that might not measure proper understanding of the task. By contrast, we evaluate the models by de-coupling these implicit signals and base our evaluation on understanding of the tasks. Lastly, we also provide a framework for breaking down claim-evidence pairs into different atomic reasoning types. While LLMs have been extensively applied in fact-checking, the question of which reasoning capabilities are needed to verify claims remains unaddressed. Thus, we are the first to propose a reasoning benchmark for claim verification. Although all three datasets used in our study can be used for fact checking, we do not explicitly cast our problem as improving fact checking performance. For instance, Hu et al. (2023b) evaluates the models’ fact checking capabilities based on models’ internal knowledge using claims only. By contrast, we provide the model with claim and evidence pairs. Therefore, our focus centers more around the models’ capabilities of understanding the evidence and implicitly reasoning to come to a conclusion in order to verify a claim. Other examples are Aly et al. (2023), Tang et al. (2024), and Strong et al. (2024), where a significant difference with our work is our choice of claims. In particular, all our claims are real-world claims and evidence. Additionally, we focus on news articles as we build on the PHEMEPlus dataset rather than on Wikipedia sources. Furthermore, we also provide a reasoning framework to facilitate the decomposition of claims into atomic reasoning types.